[Hallc_running] [New Logentry] RC Report

smithg at jlab.org smithg at jlab.org
Tue Nov 26 12:00:03 EST 2019

Logentry Text:
RC Report for Tuesday, Nov. 26

>From info gathered at the 7:45 & 8:00 meetings, plus talking to several people:

The PD says they are a little behind, but sticking to the schedule so far: A&B spin dance Saturday, hi current 5 pass to A on Sunday, beam to all halls Monday Dec. 2. Five pass energy will be 121+(1031+1031)*5 = 10.43 GeV - synchrotron losses - tweaking to optimize polarization ~ 10.4 GeV at 5 pass to Hall C. I inquired about our hot checkout list- the one item outstanding is to verify our PSS sweep pattern does not need to be changed for this experiment. I spoke to Walter about this (after JP & Brad) and he'll sign off on it. The laser platform is not enclosed, and the laser hut can't hide a person. I also talked with the PD about setting up alarms in MCC on the ramp rate and max juice. He'll get back to me tomorrow but thought this was probably no problem.

We're feeling the pressure rise as we run into the long weekend tomorrow sometime. Today Brad's getting coils placed. Then the DS spool(s) go in, get a rough alignment today hopefully, the pump gets turned on today for the entire DS beamline, and a survey check tomorrow. Walter's getting new windows for the 2 spectrometer snouts on and leak-checked today, and we're talking about only putting them on the spectrometers Monday since they would interfere with the work that needs to be done on the pivot this long weekend. Why is Monday OK to do this? Because we have to complete checkout lists in the hall on Monday anyway, always done just before we close the hall: Dave G. (US beamline), Steve L. (magnets), Brad (dets), JP (tgt), Paul Collins (final laser safety inspection), and Walter (everything else). The tentative plan is to tackle all this first thing Monday morning, and we plan to lock up at 2 or 3 PM on Monday Dec. 2. This assumes that the 3He cell goes in tomorrow, gets pu!
 mped Thursday, and the 3 target polarimeters get set up Fri-Sun. And that plan depends on getting the LSOP signed first. JP is working on this...

Also today I have to load the operational restrictions, and the myriad problems with the magnets that Brad noted in his HCLOG this morning have to be addressed. I spoke to Steve W. about this and he said he'd get on it with Mike F. The pointing survey should happen today also. The BCM tuning needs to happen today or tomorrow  (Keith Cole, EES)- Dave G. is in touch with Keith about this. The last harp checkout got done today (see ELOG https://logbooks.jlab.org/entry/3744160). The US beamline pumps are on. But DO NOT manually close VBV3H07 because there is no way to pump out the spool piece if that valve gets closed. The interlocks that normally would close it are deliberately disabled, I'm told.

Also note Mark J. HCLOG about the raster, working now at 5 mm! 

Tomorrow's 1:30 MCC meeting will be right after the 8 AM MCC meeting, because so many people want to leave early. Brad & JP asked for an RC meeting in the counting room tomorrow at 9:30 AM, when I'll hand off to Dave Mack, the next RC.


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3744210
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20191126/428cc343/attachment.html>

More information about the Hallc_running mailing list